OpenStudy (anonymous):

If xy = 5, evaluate log base 5 (x^2 *y^2). Help! This is driving me crazy!

OpenStudy (anonymous):

=2

OpenStudy (anonymous):

I think it would be 2. \[\log_{5} (x^{2} \times y^{2}) \] if xy = 5, then I think this could be substituted/simplified as \[\log_{5}(5^{2}) = 2\log_{5}(5) = 2 \times 1 = 2\]

OpenStudy (anonymous):

here xy=5 log base5(x^2y^2) =log base5(xy)2 =2logbase5(5) since log base5(5)=1 ans is 2

OpenStudy (anonymous):

Is it OK

OpenStudy (anonymous):

log base5(xy)^2 =log base5(5)^2 =2log base5(5) =2x1=2 ans

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

Thanks everyone! No questions. I knew it was an easy one but could not get it.